jhonyy9 (jhonyy9):

my name here changed to anonymous too anonymous Dear ,,Honorary Professors of Mathematics"and ,,dear Champions" ,...,to everybody ! how many ,,Super Twins" primes there may be what are twins with below property : the amount members +/- 1 result new Twin - i have got till now two twins with these property : 5-7 and 29-31 ... thank you for your all opinions,ideas !

OpenStudy (raffle_snaffle):

?

jhonyy9 (jhonyy9):

anonymous Dear ,,Honorary Professors of Mathematics"and ,,dear Champions" ,...,to everybody ! how many ,,Super Twins" primes there may be what are twins with below property : the amount members +/- 1 result new Twin - i have got till now two twins with these property : 5-7 and 29-31 ... thank you for your all opinions,ideas ! 21 days agoReport Abuse Share Question: This Question is Open 42 anonymous Best Response Medals 0 @ganeshie8 @Kainui @zepdrix @mathmale 21 days agoReport Abuse 42 anonymous Best Response Medals 0 for example 5-7 5+7=12 +/- 1 => 11 - 13 29-31 29+31=60 +/-1 => 59 - 61 there may be how many again ? 21 days agoReport Abuse 61 Bobo-i-bo Best Response Medals 1 So this must be strongly connected to the twin prime conjecture so I don't know how far we can get with this... but this is from wiki: "Every twin prime pair except (3, 5) is of the form (6n − 1, 6n + 1) for some natural number n; that is, the number between the two primes is a multiple of 6." So therefore we want twin primes, (6n−1,6n+1) to form twin primes (6m−1,6m+1) such that: (6n−1)+(6n+1)±1=6m±1 ⇒12n=6m ⇒2n=m Therefore we want to find twin primes (6n−1,6n+1) such that (12n−1,12n+1) are also twin primes... that's very basic elementary thought, I don't think i can do anymore :P 20 days agoReport Abuse 42 anonymous Best Response Medals 0 thank you this your answer but i think you dont understood it correct the above wrote property of these super twins primes - do you can getting again one 3rd twin like these two ? so than yes i wait post it here 20 days agoReport Abuse 42 anonymous Best Response Medals 0 so now i have got the 3rd twin with these above wrote properties for to be Super Twin so the 3rd Super Twin is 659 - 661 bc. 659+661 = 1320 +/- 1 = 1319 - 1321 so what are twins too how many same twins exist again there ? 20 days agoReport Abuse 42 anonymous Best Response Medals 0 woooaaawwwww !!!! here is the 4th Super Twin what is 809 - 811 so bc. 809 + 811 = 1620 +/- 1 = 1619 - 1621 what are twins too the 5th Super Twin will be 2129 - 2131 so bc. 2129 +2131 = 4260 +/- 1 = 4259 - 4261 what are twins too the 6th Super Twin is 2549 - 2551 so bc. 2549 +2551 = 5100 +/- 1 = 5099 - 5101 what are twins too 20 days agoReport Abuse 42 anonymous Best Response Medals 0 so here is the 7th Super Twin what will be 3329 - 3331 bc. 3329 + 3331 = 6660 +/- 1 = 6659 - 6661 what are twins too here is the 8th Super Twin what will be 3389 - 3391 bc. 3389 + 3391 = 6780 +/- 1 = 6779 - 6781 what are twins too here is the 9th Super Twin what will be 5849 - 5851 bc. 5849 + 5851 = 11700 +/- 1 = 11699 - 11701 what are twins too 6269 - 6271 here is the 10th Super Twin what will be 6269 - 6271 bc. 6269 + 6271 = 12540 +/- 1 = 12539 - 12541 what are twins too so how you think please can we suppose that the line of Super Twins is infinity ? 20 days agoReport Abuse 69 thomas5267 Best Response Medals 1 Given that the twin prime conjecture is still open, I don't see how we could do anything with this. Since if there are infinitely many Super Twins, then there must be infinitely twin primes, then you would have solve the twin prime conjecture and you will be famous! 20 days agoReport Abuse 42 anonymous Best Response Medals 0 @IrishBoy123 thank you very much so this wan meaning that you understand my subject about ,,Super Twins" and when i have got these 3rd,4th,...,10th why was for me like specially ? because till yestoday night i dont checked these just till 200 - 300 and how you see the 3rd is 659 - 661 ... and till the 10th 20 days agoReport Abuse 99 IrishBoy123 Best Response Medals 0 **super twins** !!!! 20 days agoReport Abuse 42 anonymous Best Response Medals 0 @IrishBoy123 so this mean that you like this ,,Super Twins" name - how i named these twins with this super propertie that the sum of members +/-1 generate,result one new twin ? thank you very very much @Directrix please your opinion about this ? - thank you in advance @Hero @satellite73 @ganeshie8 @zepdrix @mathmale 19 days agoReport Abuse qh 99 ganeshie8 Best Response Medals 0 This is very interesting as it connects two different types of rare primes : 1) Twin primes 2) Sophie Germain primes What we're looking for here is the intersection of above two sets of primes 19 days agoReport Abuse qh 99 ganeshie8 Best Response Medals 0 We don't know if there are infinitely many of either type of primes 19 days agoReport Abuse qh 99 ganeshie8 Best Response Medals 0 May I know how you're generating these "super twins" @jhonyy9 ? 19 days agoReport Abuse 42 anonymous Best Response Medals 0 @ganeshie8 there is the way how we can generate thet - just you need getting from these prims list what satisfie this propertie that the sum of members +/- 1 result one new twin ok. ? 19 days agoReport Abuse 42 anonymous Best Response Medals 0 firstly i have thought that there are just two like twins bc. i have got the 5-7 and 29-31 and i not have checked them in the greater levels - so than you see this above posted list the next are 659 - 661 so the difference between i think is to much - yes ? 19 days agoReport Abuse qh 99 ganeshie8 Best Response Medals 0 Yeah the difference between second super twin and third super twin is large 19 days agoReport Abuse 42 anonymous Best Response Medals 0 yeah @ganeshie8 than you check these super twins every member ending in 9 - 1 so 9 - 1 this is so interestingly i think - yes ? i have got this propertie of members now in this moment - do you see it ? 19 days agoReport Abuse 42 anonymous Best Response Medals 0 from this we can deducting that all tewins what has members what ending in 9 -1 there are possobolity to forme super twins - yes ? 19 days agoReport Abuse 42 anonymous Best Response Medals 0 possibility - sorry 19 days agoReport Abuse 42 anonymous Best Response Medals 0 @ganeshie8 sorry - but what you have wrote above about super twins and S. Germain prims so do you think that these super prims sometime will can be named jhony prims about my name or like somehow ? - sorry this is one my dream 19 days agoReport Abuse 42 anonymous Best Response Medals 0 @ganeshie8 one question again - how you think please ? i have some idea that i can writing a story with name,title ,,the legend of primes " how you think please will be more interestingly or not will interest nobody ? 19 days agoReport Abuse qh 99 ganeshie8 Best Response Medals 0 Very nice! so these super twins always have the form 9-1 9-1. That is, the first prime in the twin prime pair always ends in 9 and the second prime always ends in 1. 19 days agoReport Abuse 42 anonymous Best Response Medals 0 @TheSmartOne here can reading about super twins - hope is understandably - wait your opinion about these - thank you 13 days agoReport Abuse 99 TheSmartOne Best Response Medals 0 * very interesting 13 days agoReport Abuse 42 anonymous Best Response Medals 0 ty 13 days agoReport Abuse 42 anonymous Best Response Medals 0 so and what about star twins after you know these about super t 13 days agoReport Abuse 42 anonymous Best Response Medals 0 super twins 13 days agoReport Abuse 42 anonymous Best Response Medals 0 ganeshie8 Honorary Professor of Mathematics Best Response Medals 0 This is very interesting as it connects two different types of rare primes : 1) Twin primes 2) Sophie Germain primes What we're looking for here is the intersection of above two sets of primes ganeshie8 Honorary Professor of Mathematics Best Response Medals 0 We don't know if there are infinitely many of either type of primes 5 days agoReport Abuse qh 99 ganeshie8 Honorary Professor of Mathematics Best Response Medals 0 May I know how you're generating these "super twins" @jhonyy9 ? @TheSmartOne 13 days agoReport Abuse 42 anonymous Best Response Medals 0 @jim_thompson5910 @ikram002p here you can reading about super twins - opinions please - thank you very much 11 days agoReport Abuse Type your reply

OpenStudy (zzr0ck3r):

is this a joke?

OpenStudy (zzr0ck3r):

Polignac's conjecture from 1849 states that for every positive even natural number k, there are infinitely many consecutive prime pairs p and p′ such that p′ − p = k (i.e. there are infinitely many prime gaps of size k). The case k = 2 is the twin prime conjecture.